# Chiral Gapless Spin Liquid in Hyperbolic Space

### v2

@@ Introduction.
where $\hat{P}$ is the projector onto states with negative energy; $A$, $B$, and $C$ are three sectors in counterclockwise order that evenly partition the hyperbolic flake and do not include boundary sites (see [33], Sec. E). We find the LCM to vanish everywhere in the dimer phase $A$, but to be finite-valued in the gapless phase for either one of the vortex-filled or vortex-free sectors (cf. Fig. 4(a)). We observe that the value of the Chern number is not quantized. This is not inconsistent with the expected phenomenology, since the Chern number of a topological band is only quantized if the Fermi level resides within a band gap. Furthermore, the local Chern marker calculation might be affected due to finite size or edge effects in the hyperbolic plane.
Refer to caption Refer to caption $t=0$Refer to caption $t=2000$Refer to caption $t=4000$Refer to caption $t=6000$(a)(b)
Figure 4: (a) Phase diagram of the local Chern marker (LCM, cf. Eq. (6)) in the plane $J_{x}+J_{y}+J_{z}=1$ for a hyperbolic droplet with 1575 sites. The LCM is calculated from (3) including all states with negative energy. It is finite but not quantized in the gapless $B$ phase, and vanishes in the $A$ phases. (b) Time evolution under (3) of a wave packet initially localized at the edge in the gapless chiral phase and the uniform flux sector. The center of the packet propagates unidirectionally along the boundary. The area of each circle is proportional to the probability of the excitation to be located on the corresponding site.
In order to investigate the chiral character of the gapless phase $B$ from a different angle without reference to band theory, we conduct a numerical propagation experiment in the Majorana representation of (1). We apply the time evolution of (3) in the vortex-free configuration to an initial state with energy expectation value $\langle H\rangle=0$, which represents an excitation localized at the edge of a droplet (Fig. 4(b), see [33], Sec. E). After $2000$, $4000$, and $6000$ time steps we observe that the excitation has moved unidirectionally along the edge in counterclockwise direction. This indicates the chiral character of the eigenstates in the vortex-free sector. As expected, the direction of propagation is reversed for the vortex-filled sector.
@@
As a future endeavor, it will be important to study the behavior of the spin-spin correlation function of the gapless chiral spin liquid and obtain the dynamic exponent. This would provide insights into the precise nature of this ground state. Furthermore, one should study its stability against different types of perturbations. From here, utilizing a projective symmetry group (PSG) classification of spin liquids towards characterizing the microscopic symmetries of the underlying quantum orders in hyperbolic space presents itself as a natural thread of investigation. This is because the PSG may enable us to identify the symmetries protecting the gapless chiral excitations, and to study their intertwining with a potentially non-Abelian Bloch translation group profile.
@@
Upon the finalization of this manuscript, we became aware of an independent simultaneous work addressing hyperbolic Kitaev spin liquids on the $\{8,3\}$ lattice with and without the presence of an external magnetic field [39].
Acknowledgement.
We thank A. Stegmaier, I. Boettcher, P. M. Lenggenhager, S. Dey, T. Bzdušek, and J. Maciejko for discussions. The work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 258499086 - SFB 1170 and through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat Project-ID 390858490 - EXC 2147 as well as by the National Science Foundation through NSF-PHY-2210452.
@@ Acknowledgement.
Supplemental Material A Ground state flux sector
@@
The Kitaev Hamiltonian (Eq. (1) of the main text) commutes with the plaquette operators $\hat{W}_{p}$. Therefore, the Hilbert space can be decomposed into the eigenspaces of $\hat{W}_{p}$, which we call flux sectors. They are labeled by the eigenvalues of each plaquette operator $W_{p}=\pm\mathrm{i}$. The eigenvalues $W_{p}$ can be expressed in terms of the eigenvalues of the link operators, $u_{jk}=\pm 1$ (see Eq. (4) of the main text). A given flux configuration on the plaquettes of a lattice can arise from multiple distinct link configurations. If all $u_{jk}=+1$, one finds $W_{p}=-\mathrm{i}\;\forall p$, i.e., each plaquette carries a flux of $-\frac{\pi}{2}$. We call this factor the vortex-free sector. If all $u_{jk}=-1$, the plaquette eigenvalues are $W_{P}=+\mathrm{i}\;\forall p$. Then each plaquette carries a flux of $\frac{\pi}{2}$. This is the vortex-filled sector. Both map onto each other under time reversal, and form a Kramers pair.
@@
Now, the task is to identify the flux sector in which the ground state is located. For bipartite lattice Lieb’s theorem [32] can be employed. Regarding lattices consisting of regular $p$-gons it states that if the graph is reflection symmetric with respect to an axis that does not intersect with any lattice site, the ground state always lies in either the sector where all plaquettes have flux 00 or flux $\pi$. If $p=2\mod 4$ (as for the Kitaev Honeycomb Model), the ground state lies in the zero-flux sector, while for $p=0\mod 4$, it is located in the $\pi$-flux sector. Lieb’s theorem is not restricted to Euclidean lattices. However, it is not applicable to lattices with odd loops, since they do not possess the necessary reflection symmetry.
Refer to caption (a)(b)
Figure A.1: Energy cost of creating vortices and aggregating them in a cluster on a droplet with 7569 sites. While the energy cost per vertex decreases with the number of vortices (a), the total energy cost increases with each vortex (b). The energy cost per vertex is positive with respect to the ground state, which means that the configuration with vortices is energetically less favorable than the vortex-free state. We therefore conjecture that the ground state lies in the uniform flux sector.
Therefore, we resort to numerical means in order to identify the ground state flux sector. As a first numerical experiment, we explicitly sampled every distinct flux configuration of a droplet with 40 sites at the isotropic point $J_{x}=J_{y}=J_{z}$ and found that the energy is minimized only by the vortex-free and vortex-filled configurations. These correspond to a flux of ±$\pm\frac{\pi}{2}$ through each plaquette. As a second step we recorded the energy cost of nucleating a pair of isolated vortices in adjacent plaquettes and pulling them apart (see Fig. 2 of the main text). The results indicate, that there is an attractive force between vortices, which suggests that the formation of clusters is energetically favored. Thus, we investigated this configuration further: In Fig. A.1 we show the energy cost of adding a single vortex to a cluster and the total energy of the cluster above the flux-free sector. We find that the energy cost per vortex is always larger than zero. While it decreases with cluster size, the total energy of the cluster increases. Thus, although these numerical studies are limited to finite-sized samples, all results indicate that the ground state is formed by the Kramers doublet of the flux ±$\pm\pi$ sectors.
@@ B.1 Continued Fraction Method
The general idea of the continued fraction method [40, 41, 31] for calculating the bulk density of states (DOS) is to compute one element of the resolvent in the local basis, which is equivalent to compute the moments of the local normalized DOS. Let a physical system be described by a tight-binding Hamiltonian $\hat{H}$, and define its resolvent as
@@
for a complex number $z\in\mathbb{C}$ in the upper half of the complex plane ($\Im[z]>0$). Furthermore, let {|$\{\ket{j}\}$ represent the local basis of the single-particle Hilbert space. Then, the local DOS at site $j$ is given by
$\rho_{j}(E)=-\frac{1}{\pi}\,\lim_{\eta\rightarrow 0}\,\Im\left[\braket{j}{\hat{G}(E+\mathrm{i}\eta)}{j}\right].$ (B.2)
In order to estimate the bulk DOS of the Majorana Hamiltonian (3) in the hyperbolic plane, we consider the vertex centered $\{9,3\}$ lattice, and calculate the local DOS for the center vertex with label $j=0$. Then, the resolvent can written in terms of the moments $\mu_{0,r}$ of the Hamiltonian [40]
$G_{00}(z)=\frac{1}{z}\left(1+\sum_{r}\frac{\mu_{0,r}}{z^{3}}\right)\,,$ (B.3)
@@
We can interpret this representation of the resolvent as a sum over all loops through the lattice starting at site 00 and ending at site 00 with lengths $r=0,1,2,\cdots$. Since the moments of the Hamiltonian $\mu_{0,r}$ at site 00 are also the moments of the local DOS at site 00, finding $\mu_{0,r}$ means finding the DOS. In principle, it is necessary to calculate all moments.
@@
The coefficients $a_{n},b_{n}\in\mathbb{R}$ in the tridiagonal Hamiltonian representation are calculated using the following recursion: At the start set |$\ket{\widetilde{0}}=\ket{0}$, and $b_{0}=0$. Then we can recursively calculate the following sequence:
@@
In principle, this iteration can be repeated indefinitely, and we would obtain all moments $\mu_{0,r}$ of the local DOS. In practice for a finite system, it is only possible to determine a part of the moments exactly. When the lengths of the loops described by equation (B.4) reach the edge of the sample, the moments will deviate from the moments of the bulk DOS. This is also the iteration order $N$, after which we terminate the continued fraction (B.6).
There are different approaches to terminate the continues fraction [42]. In the simplest case, we set $g_{n}(z)\approx g_{\infty}(z)$ for $n>N$, where $g_{\infty}$ is determined by the equation
@@
In principle, we can set $a_{\infty}=a_{N}$ and $b_{\infty}=b_{N}$. However, we obtain better results by averaging over a suitable set of coefficients before the cutoff at index $N$ (see Fig. 3(a)). From the termination term $g_{\infty}(z)$, it becomes clear that the $a_{\infty}$ can be interpreted as the center of the local DOS, while $b_{\infty}$ corresponds to the bandwidth. Note that for our system we find $a_{n}\equiv 0$, as the spectrum is symmetric with respect to $E=0$.
This termination procedure works reasonably well for systems without a band gap. If there is a band gap, the termination should be chosen differently. In this case, the coefficients $b_{n}$ do not converge to a single value, but alternate between two attractors for large $n$ (see Fig. 3(a) of the main text as well as Ref. [42]). Then, the results become more precise when this behavior is reflected in the termination. For $a_{n}=0$, we distinguish $b_{\infty}^{\text{odd}}=\lim_{k\rightarrow\infty}b_{2k+1}$ and $b_{\infty}^{\text{even}}=\lim_{k\rightarrow\infty}b_{2k}$ and set $g_{n}(z)\approx g_{\infty}(z)$ for $n>N$ with $N$ even, such that
@@ B.2 Comparison with Exact Diagonalization and Hyperbolic Band Theory results
(a) $J_{x}=1/6$, $J_{y}=J_{z}=5/12$Refer to caption (b) $J_{x}=J_{y}=J_{z}=1/3$Refer to caption (c) $J_{x}=2/3$, $J_{y}=J_{z}=1/6$Refer to caption (d) $J_{x}=1/2$, $J_{y}=J_{z}=1/4$Refer to caption
@@
We compare the results obtained by the continued fraction (CF) method with results from Exact Diagonalization (ED) and Hyperbolic Band Theory (HBT) methods. The ED results are obtained by diagonalizing the Majorana Hamiltonian (3) for a droplet configuration with 33 3883338833\,38833 388 sites and open boundary conditions (OBC). From the spectrum, we deduce the DOS
@@
In order to calculate the spectrum from Hyperbolic Band Theory, we employ the following trick: Since we are only interested in the eigenvalues, we can apply any similarity transform to the Hamiltonian as long as the flux of ±$\pm\pi/2$ through a plaquette is invariant. With such a mapping, the Majorana Hamiltonian can be transformed into a single-particle Hamiltonian with a constant magnetic field. This problem was already solved for periodic boundary conditions in Reference [18]. For these results we use a compactified hyperbolic graph with $108$ sites. We calculate its hyperbolic Bloch form and sample the momenta randomly.
Figure B.1 compares the results of the three methods. We observe that the DOS computed by the three methods agree qualitatively. Regarding the pronounced local minima in the DOS, for example in the $J_{x}=J_{y}=J_{z}=1/3$ case, all three methods predict their presence. But there are mismatches in the exact location and the depth of the minima. Most importantly, however, all three methods agree that there are no gaps in the spectrum within the chiral phase.
The deviations arise since all three methods have their drawbacks: The ED methods works only for sufficiently small droplet sizes. Moreover, the categorization of sites belonging to the bulk or to the edge by their position in the Poincaré disk depends on the choice of the cutoff: If it is too small, important contributions to the bulk DOS may be missing, if the cutoff is too large, the DOS picks up edge effects. Furthermore, the results may depend on the binning procedure. The results using Hyperbolic Band Theory on a compactified graph can also suffer from finite size effects. Depending on the size of the unit cell, contributions to the DOS coming from higher-order representations of the translational group may be missing. The CF method is applicable to much larger droplet sizes, since the method requires only calculating matrix-vector products. Still, only a certain number of moments will be exact, before the continued fraction must be terminated. This can lead to convergence problems, especially close to sharp peaks or dips in the DOS.
We conclude that the combination of all three methods is suitable to predict qualitative features of the DOS of the Majorana Hamiltonian on the $\{9,3\}$ lattice. In the main text, we use these results to check whether the ground state Hamiltonian is gapped.
Supplemental Material C Toric Code Limit and perturbation theory in the gapped phase

### v4

@@ Introduction.
where $\hat{P}$ is the projector onto states with negative energy; $A$, $B$, and $C$ denote three sectors in counterclockwise order that evenly partition the hyperbolic flake not including boundary sites (see [36], Sec. E). We find the LCM to vanish everywhere in the dimer phase $A$, but to be finite valued in the gapless phase for either one of the ±$\pm\pi/2$ flux sectors (cf. Fig. 4(a)). We observe that the value of the LCM is not quantized. This is not inconsistent with the expected phenomenology, since the Chern number of a topological band is only quantized if the Fermi level resides within a band gap. Furthermore, the LCM calculation might be affected due to finite-size or edge effects in the hyperbolic plane.
In order to investigate the chiral character of the gapless phase $B$ from a different angle, we conduct a numerical propagation experiment in the Majorana representation of (1). We apply the time evolution of (2) in the $+\pi/2$ flux configuration to an initial state with energy expectation value $\langle H\rangle=0$, which represents an excitation localized at the edge of a droplet (Fig. 4(b), see [36], Sec. F). After $800$, $1600$, and $2400$ time steps we observe that the excitation has moved unidirectionally along the edge. As expected, the direction of propagation is reversed for the $-\pi/2$ flux sector.
Perturbative magnetic field.
Under the application of a small magnetic field, the Hamiltonian becomes $\hat{\widetilde{H}}=\hat{H}+\hat{H^{\prime}}$ with
$\hat{H}^{\prime}=-\kappa\sum_{\begin{subarray}{c}(j,k,l)_{\alpha\gamma}\\ \beta\neq\alpha,\beta\neq\gamma\end{subarray}}\hat{\sigma}_{j}^{\alpha}\,\hat{\sigma}_{k}^{\beta}\,\hat{\sigma}_{l}^{\gamma}.$ (5)
The sum runs over all three-tuples of sites $(j,k,l)_{\alpha\gamma}$, such that sites $j$ and $k$ are connected anticlockwise around a plaquette by an $\alpha$-type bond, and $k$ and $l$ by a $\gamma$-type bond. $\beta$ is then given by the third bond flavor. This simple form is sustained by the regular coloring characteristic for the $\{9,3\}$ lattice. In the Majorana language the Hamiltonian takes the form
$\hat{H}^{\prime}=\mathrm{i}\,\frac{\kappa}{2}\sum_{(j,k,l)}\hat{u}_{jk}\hat{u}_{kl}\hat{c}_{j}\hat{c}_{l},$ (6)
which corresponds to a second-nearest neighbor hopping term for the Majorana fermions.
We set $J_{x}=J_{y}=J_{z}=\frac{1}{3}$ and analyze the model for $\kappa\neq 0$. As the magnetic term breaks time-reversal symmetry, the ±$\pm\pi/2$ flux sectors are no longer degenerate. For $\kappa>0$, the ground state is given by the uniform $-\pi/2$ configuration, while the $+\pi/2$ configuration constitutes the ground state for $\kappa<0$. We compute the fermion gap for both sectors (Fig. 2 (c)), and find that a gap opens for arbitrarily small $\kappa$.
@@
As a future endeavor, it will be important to study the behavior of the spin-spin correlation function of the gapless chiral spin liquid and obtain the dynamic exponent. This would provide insights into the precise nature of this ground state. From here, utilizing a projective symmetry group (PSG) classification of spin liquids towards characterizing the microscopic symmetries of the underlying quantum orders in hyperbolic space presents itself as a natural thread of investigation. This is because the PSG may enable us to identify the symmetries protecting the gapless chiral excitations, and to study their intertwining with a potentially non-Abelian Bloch translation group profile.
@@
Recently, we became aware of an independent simultaneous work addressing hyperbolic Kitaev spin liquids on the $\{8,3\}$ lattice with and without the presence of an external magnetic field [45].
Acknowledgements.
We thank A. Stegmaier, I. Boettcher, P. M. Lenggenhager, S. Dey, T. Bzdušek, and J. Maciejko for discussions, as well as R. Vogeler for providing a large PBC cluster of the $\{9,3\}$ hyperbolic lattice. The work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 258499086 - SFB 1170 and through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat Project-ID 390858490 - EXC 2147 as well as by the National Science Foundation through NSF-PHY-2210452.
Author contributions.
R.T. initiated and supervised the project. F.D., T.H., and R.T. designed the model system. F.D. and T.H. carried out the calculations with discursive and calculation input from J.V., R.M., A.M., M.G., and Y.I.. All authors contributed to writing the manuscript.
@@ Acknowledgements.
Supplemental Material A Ground state flux sector and excitations
@@
The Kitaev Hamiltonian (Eq. (1) of the main text) commutes with the plaquette operators $\hat{W}_{p}$. Therefore, the Hilbert space can be decomposed into the eigenspaces of $\hat{W}_{p}$, which we call flux sectors. They are labeled by the eigenvalues of each plaquette operator $W_{p}=\pm\mathrm{i}$. The eigenvalues $W_{p}$ can be expressed in terms of the eigenvalues of the link operators, $u_{jk}=\pm 1$. A given flux configuration on the plaquettes of a lattice can arise from multiple distinct link configurations. If all $u_{jk}=+1$, one finds $W_{p}=-\mathrm{i}\;\forall p$, i.e., each plaquette carries a flux of $-\frac{\pi}{2}$. Accordingly, we call this sector the uniform $-\pi/2$ flux sector. If all $u_{jk}=-1$, the plaquette eigenvalues are $W_{p}=+\mathrm{i}\;\forall p$. Then each plaquette carries a flux of $\frac{\pi}{2}$. Both map onto each other under time reversal, and form a Kramers pair.
@@
We wish to identify the flux sector in which the ground state is located. For bipartite lattices Lieb’s theorem [35] can be employed. Regarding lattices consisting of regular $p$-gons it states that if the graph is reflection symmetric with respect to an axis that does not intersect with any lattice site, the ground state always lies in either the sector where all plaquettes have flux 00 or flux $\pi$. If $p=2\mod 4$ (as for the Kitaev Honeycomb Model), the ground state lies in the uniform zero-flux sector, while for $p=0\mod 4$, it is located in the uniform $\pi$-flux sector. Lieb’s theorem is not restricted to Euclidean lattices. However, it is not applicable to lattices with odd loops, since they do not possess the necessary reflection symmetry. Nevertheless, we conjecture that the ground states lie in the uniform flux sectors, and resort to numerical means in order to validate this conjecture and to characterize the corresponding excitations.
Refer to caption Refer to caption (a)(b)(c)
Figure A.1: (a) Energy cost of creating a pair of spatially separated visons in a cluster with 11878 sites for the parameters $J_{x}=J_{y}=J_{z}=\frac{1}{3}$. The energy is increasing with the distance between the visons, hinting at an attractive force between the excitations. (b) Energy cost of creating visons and aggregating them in a cluster on a droplet with 7569 sites for the parameters $J_{x}=J_{y}=J_{z}=\frac{1}{3}$. The total energy cost increases approximately linearly with the number of visons. (c) Energy cost of adding visons to the cluster per vison for the same configuration as in (b). The energy cost per vison is positive with respect to the ground state, which means that the configuration with vortices is energetically less favored than the conjectured uniform flux ground state. Note that all results may exhibit finite-size effects, such that the quantitative findings do not necessarily transfer to the thermodynamic limit.
As a first basic numerical experiment, we explicitly sample every distinct flux configuration of a droplet with 40 sites at the isotropic point $J_{x}=J_{y}=J_{z}$ and find that the energy is minimized only by the uniform ±$\pm\frac{\pi}{2}$ flux configurations.
Furthermore, we analyze the energy increase due to the creation of vortex excitations (visons) over the assumed ground states in finite-size OBC clusters. For this, we consider the chiral phase $B$ including the isotropic coupling point, where $J_{x}=J_{y}=J_{z}$. (For an analysis of the dimer phases $A_{\alpha}$, where the dimer limit can be compared to analytical calculations, refer to Section C of the Supplemental Material.) We compute the energy cost of nucleating a pair of isolated visons in adjacent plaquettes and pulling them apart. The result is shown in Figure A.1 (a). Although we acknowledge that these results will quantitatively depend on the cluster size (as it is the case in the $A$ phases (c.f. Sec. C.2)), they indicate qualitatively that there is an attractive force between vortices. This suggests that the formation of vison clusters is energetically favored.
Thus, the configuration of vison clusters is investigated further: In Fig. A.1 (b), (c) we show the energy cost of adding a single vison to a cluster as well as the total energy of the cluster above the uniform flux sector. We find that the energy cost per vison is always larger than zero and that the total energy of the cluster increases linearly with the number of visons. This result is reasonable, as we expect the energy of the cluster to be proportional to the length of the cluster’s boundary, which, in turn, is proportional to the volume of the cluster. Thus, although these numerical studies are limited to finite-sized samples, all results indicate that the ground state is formed by the Kramers doublet of the flux ±$\pm\pi/2$ sectors.
@@ B.1 Continued Fraction Method
The general idea of the continued fraction (CF) method [46, 47, 32] for calculating the bulk density of states (DOS) is to compute one element of the resolvent in the local basis, which is equivalent to compute the moments of the local normalized DOS. Let a physical system be described by a tight-binding Hamiltonian $\hat{H}$, and define its resolvent as
@@
for a complex number $z\in\mathbb{C}$ in the upper half of the complex plane ($\mathrm{Im}[z]>0$). Furthermore, let {|$\{\ket{j}\}$ represent the local basis of the single-particle Hilbert space. Then, the local DOS at site $j$ is given by
$\rho_{j}(E)=-\frac{1}{\pi}\,\lim_{\eta\rightarrow 0}\,\mathrm{Im}\left[\braket{j}{\hat{G}(E+\mathrm{i}\eta)}{j}\right].$ (B.2)
In order to estimate the bulk DOS of the Majorana Hamiltonian (2) in the hyperbolic plane, we consider the vertex-centered $\{9,3\}$ lattice, and calculate the local DOS for the central vertex with label $j=0$. Then, the resolvent can be written in terms of the moments $\mu_{0,r}$ of the Hamiltonian [46]
$G_{00}(z)=\frac{1}{z}\left(1+\sum_{r}\frac{\mu_{0,r}}{z^{r}}\right)\,,$ (B.3)
@@
We can interpret this representation of the resolvent as a sum over all loops through the lattice starting and ending at site 00 with lengths $r=0,1,2,\cdots$. Since the moments of the Hamiltonian $\mu_{0,r}$ at site 00 are also the moments of the local DOS at site 00, finding $\mu_{0,r}$ means finding the DOS. In principle, it is necessary to calculate all moments.
@@
The coefficients $a_{n},b_{n}\in\mathbb{R}$ in the tridiagonal Hamiltonian representation are calculated using the following recursion: First, we set |$\ket{\widetilde{0}}=\ket{0}$, and $b_{0}=0$. Then we can recursively calculate the following sequence:
@@
In principle, this iteration can be repeated indefinitely, and we would obtain all moments $\mu_{0,r}$ of the local DOS. In practice, it is only possible to determine a part of the moments exactly for a finite system. As the length of the loops described by equation (B.4) increases, they will reach the edge of the sample and the moments will deviate from the moments of the bulk DOS. This is also the iteration order $N$, after which we terminate the continued fraction (B.6).
For periodic gapless structures, the coefficients $a_{n}$ $b_{n}$ converge slowly towards their respective asymptotic values $a_{\infty}$, $b_{\infty}$, with oscillations associated with van Hove singularities in the spectrum. In the ideal case, the coefficients are already sufficiently converged at iteration order $N$, which is the case, for example, for gapless regular hyperbolic tilings [37].
Then, we set $g_{n}(z)\approx g_{\infty}(z)$ for $n>N$, where $g_{\infty}$ is determined by the self-consistent equation
@@
In the simplest case, where the coefficients are converged, we identify $a_{\infty}=a_{N}$ and $b_{\infty}=b_{N}$.
For our model, the convergence is much slower than for regular hyperbolic tiling. Thus, we obtain better results by setting $b_{\infty}$ to the average over a suitable set of coefficients before the cutoff at index $N$ (see Fig. 2(a)). From the termination term $g_{\infty}(z)$, it becomes clear that the $a_{\infty}$ can be interpreted as the center of the local DOS, while $b_{\infty}$ corresponds to the bandwidth. Note that for our system we find $a_{n}\equiv 0$, as the spectrum is symmetric with respect to $E=0$.
This termination procedure works reasonably well as long as the DOS is not gapped. In the presence of a gap, the termination should be chosen differently [48]. In this case, the coefficients $b_{n}$ do not converge to a single value, but alternate between two attractors for large $n$ (see Fig. 2(a) of the main text as well as Ref. [48]). Then, the results become more precise when this behavior is reflected in the termination. For $a_{n}=0$, we distinguish $b_{\infty}^{\text{odd}}=\lim_{k\rightarrow\infty}b_{2k+1}$ and $b_{\infty}^{\text{even}}=\lim_{k\rightarrow\infty}b_{2k}$ and set $g_{n}(z)\approx g_{\infty}(z)$ for $n>N$ with $N$ even, such that
@@ B.2 Comparison with Exact Diagonalization and Hyperbolic Band Theory results
(a) $J_{x}=1/6$, $J_{y}=J_{z}=5/12$Refer to caption (b) $J_{x}=J_{y}=J_{z}=1/3$Refer to caption (c) $J_{x}=1/2$, $J_{y}=J_{z}=1/4$Refer to caption (d) $J_{x}=2/3$, $J_{y}=J_{z}=1/6$Refer to caption
@@
We compare the results obtained by the continued fraction (CF) method with results from Exact Diagonalization (ED) and Hyperbolic Band Theory (HBT) methods. The ED results are obtained by diagonalizing the Majorana Hamiltonian (2) for a droplet configuration with 33 3883338833\,38833 388 sites and open boundary conditions (OBC) as well as for periodic boundary conditions (PBC) in the case $J_{x}=J_{y}=J_{z}=\frac{1}{3}$. From the spectrum, we deduce the DOS
@@
In order to calculate the spectrum from HBT, we employ the following trick: Since we are only interested in the eigenvalues, we can apply any similarity transform to the Hamiltonian as long as the flux of ±$\pm\pi/2$ through a plaquette is invariant. With such a mapping, the Majorana Hamiltonian can be transformed into a single-particle Hamiltonian with a constant magnetic field. This problem was already solved for periodic boundary conditions in Reference [18]. For these results we use a compactified hyperbolic graph with $108$ sites. We calculate its hyperbolic Bloch Hamiltonian and sample the momenta randomly.
Figure B.1 compares the resulting DOS of the three methods. We observe that the results agree qualitatively. For example, all methods predict the presence of the pronounced local minima in the $J_{x}=J_{y}=J_{z}=1/3$ case. But there are mismatches in the exact location and the depth of the minima. Most importantly, however, all methods agree that there are no gaps in the spectrum within the chiral phase.
The deviations arise since all three methods have their drawbacks: The ED methods works only for sufficiently small droplet sizes. Moreover, the categorization of sites belonging to the bulk or to the edge by their position in the Poincaré disk depends on the choice of the cutoff: If it is too small, important contributions to the bulk DOS may be missing, if the cutoff is too large, the DOS picks up edge effects. Furthermore, the results may depend on the binning procedure. These issues seem to be mitigated for the PBC cluster. However, such a PBC calculation was only performed for the isotropic point $J_{x}=J_{y}=J_{z}$ where the coloring of the graph does not play a role. The results using Hyperbolic Band Theory on a compactified graph can also suffer from finite-size effects. Depending on the size of the unit cell, contributions to the DOS coming from higher-order representations of the translational group may be missing. The CF method is applicable to much larger droplet sizes, since the method requires only calculating matrix-vector products. Still, only a certain number of moments will be exact, before the continued fraction must be terminated. This can lead to convergence problems, especially close to sharp peaks or dips in the DOS. As we see in Fig. B.1 (b), the CF results exhibit ripple artifacts oscillating around the PBC ED result.
Despite the drawbacks of the individual methods, we conclude that the combination of all three methods is suitable to predict qualitative features of the DOS of the Majorana Hamiltonian on the $\{9,3\}$ lattice.
Supplemental Material C Dimer limit and perturbation theory in the gapped phase
C.1 Derivation of the effective Hamiltonian for vison excitations in the dimer limit